# A Consistent Cosmic Shear Analysis in Harmonic and Real Space

### v1

@@ Abstract
Recent cosmic shear analyses have exhibited inconsistencies of up to $1\sigma$ between the inferred cosmological parameters when analyzing summary statistics in real space versus harmonic space. In this paper, we demonstrate the consistent measurement and analysis of cosmic shear two-point functions in harmonic and real space using the $i$Master algorithm. This algorithm provides a consistent prescription to model the survey window effects and scale cuts in both real space (due to observational systematics) and harmonic space (due to model limitations), resulting in a consistent estimation of the cosmic shear power spectrum from both harmonic and real space estimators. We show that the $i$Master algorithm gives consistent results using measurements from the HSC Y1 mock shape catalogs in both real and harmonic space, resulting in consistent inferences of $S_{8}=\sigma_{8}(\Omega_{m}/0.3)^{0.5}$. This method provides an unbiased estimate of the cosmic shear power spectrum, and $S_{8}$ inference that has a correlation coefficient of 0.997 between analyses using measurements in real space and harmonic space. We observe the mean difference between the two inferred $S_{8}$ values to be 0.0004, far below the observed difference of 0.042 for the published HSC Y1 analyses and well below the statistical uncertainties. While the notation employed in this paper is specific to photometric galaxy surveys, the methods are equally applicable and can be extended to spectroscopic galaxy surveys, intensity mapping, and CMB surveys.
@@
pubyear: 2023pagerange: A Consistent Cosmic Shear Analysis in Harmonic and Real SpaceB
@@ 1 Introduction
Once the data is acquired and processed into shear maps or shear catalogs, the next step is to compute the two-point summary statistics. One question that we need to address at this stage is whether to carry out the cosmic shear analysis in real space or harmonic space; each space has its own advantages and disadvantages, and we introduce the trade-offs in Sec. 2.3. Analysis in real space corresponds to measuring the two-point correlation functions of the galaxy shear estimates as a function of angular separation (e.g., as in Li et al., 2023; Secco et al., 2022), and analysis in harmonic space corresponds to measuring the cosmic shear power spectrum as a function of multipole moments (e.g., as in Dalal et al., 2023; Doux et al., 2022). One major challenge is to account for the survey window. The survey window is a function that captures the effects of the survey geometry and other observational effects that affect the selection of galaxies included in our samples. The survey window complicates the power spectrum analysis as the survey mask couples Fourier modes, which would otherwise be independent. This leads to mode coupling in the power spectrum, which is commonly addressed using pseudo-$C_{\ell}$ estimators (e.g., Wandelt et al., 2001; Hikage et al., 2011; Alonso et al., 2019)
@@
Since the coherent distortion of galaxy shapes is caused by a scalar gravitational potential of the matter distribution, we can further separate the lensing signal into curl-free ($E$ mode) and divergence-free ($B$ mode) components. The survey window also introduces leakage of $E$ modes into the $B$ mode component of a given map, and vice versa, an effect that needs to be modeled and corrected using detailed knowledge of the window. The modeling of the survey window in harmonic space was first addressed in Hivon et al. (2002), who introduced the Master algorithm to deconvolve the effects of the survey window to reconstruct the power spectrum of the underlying fields for Cosmological Background Microwave (CMB) analysis (Wandelt et al., 2001; Brown et al., 2005; Lewis et al., 2001). Alonso et al. (2019) introduced the now standard NaMaster software with an implementation of the Master algorithm for spin-0 and spin-2 fields that is also used for galaxy clustering and cosmic shear analyses. The survey window also affects the two-point correlation functions in real space, although this effect is typically ignored since the commonly-used estimator removes most of the window effects (see appendix D of Singh et al. (2020) for demonstration of residual percent-level window effects on correlation functions).
@@
After measuring the summary statistics, we apply scale cuts on the measurements to restrict our analysis to scales where the model is reliable (normally done in harmonic space where the models are defined) and to remove scales contaminated by observational systematics (normally done in real space to account for physical and observational effects tied to angular separation) before performing cosmological parameter inference. We can typically model a limited range of scales that is smaller than the range of scales probed by the current data. Thus, we need to apply scale cuts carefully to extract as much information as possible while avoiding biases from applying the model outside its range of validity. Failure to account for the scale cut that was implemented in one space when modeling the other space can lead to using different sets of information in real versus harmonic space analysis, even though the underlying field is the same. For example, Hamana et al. (2020) and Doux et al. (2021) show how such a disparity in approach produces apparent inconsistency between the analyses in real space and harmonic space when performing cosmological parameter inference. Crucially, if we truly believe certain scales cannot be well modeled, removing them regardless of the space being used for the analysis is important.
@@
There have been many efforts to resolve this inconsistency in analysis between the two spaces. Schneider et al. (2002) introduced the band power spectra that used the aperture mass dispersion $\langle M^{2}_{\text{ap}}(\theta)\rangle$ and $\langle M^{2}_{\times}(\theta)\rangle$ (Schneider et al., 1998) measured from a linear combination of the two-point correlation functions weighted by specific window functions to perform the $E/B$ mode decomposition. However, calculating the aperture mass dispersion requires estimation of the shear correlation functions down to zero angular separation, which is not possible in practice. To remedy this, they applied apodization in the window function that truncates the integral over all angular distances and reduces the ringing effect caused by the limited range of the two-point correlation functions. Schneider et al. (2010) introduced the Complete Orthogonal Sets of E- / B-Mode Integrals (COSEBI) method to measure pure $E$ and $B$-mode cosmic shear power spectra from the two-point correlation functions over a finite interval $\theta_{\text{min}}\leq\theta\leq\theta_{\text{max}}$. The method, similar to Schneider et al. (2002), uses a linear combination of binned measured two-point correlation functions weighted by specific window functions in the angular scale. Becker & Rozo (2016) introduced a similar band-power estimator for cosmic shear power spectra from linear combinations of binned two-point correlation functions. Asgari et al. (2021) used three different two-point statistics (band power spectra, COSEBIs, and two-point correlation functions) on the KiDS data set and compared the measured values of the parameter $S_{8}\equiv\sigma_{8}(\Omega_{m}/0.3)^{0.5}$, where $\sigma_{8}$ is the standard deviation of matter density fluctuations in spheres of radius $8~{}\text{Mpc}/h$ (here $h\equiv H_{0}/100)$ and $\Omega_{m}$ is the fraction of matter density in the Universe with respect to the critical density. Comparing the inferred $S_{8}$ from each method provided a useful internal consistency check between the methods; Asgari et al. (2021) found a $0.4\sigma$ difference in $S_{8}$ between the COSEBIs and the two-point correlation functions analyses. This offset can arise because the different two-point statistics have different sensitivities to $\ell$ scales, even for the same data set.
@@
Singh (2021) further improved the Master algorithm (to produce $i$Master) and generalized the method to account for the scale cuts used in the analysis, making the power spectra and two-point correlation function consistent. In this work, we present the first application of the $i$Master algorithm to survey mock catalogs with complex geometry. We measure the cosmic shear two-point functions in harmonic and real space in the HSC Y1 mock shape catalogs (Shirasaki et al., 2019), using the $i$Master method to estimate the power spectrum from both harmonic and real space estimators consistently, accounting for both the survey window and the scale cuts used in the analysis. We quantify the consistency between the two analyses by measuring $\delta S_{8}=S_{8|R\to F}-S_{8|F}$, where $S_{8|R\to F}$ is the inferred $S_{8}$ value going from the real space estimator to the harmonic space estimator, and $S_{8|F}$ using the harmonic space estimator directly.
@@
This paper is organized as follows: In Sec. 2, we provide an overview of cosmic shear analysis. In Sec. 3, we describe how two-point estimators are measured in harmonic and real space and outline steps to make a consistent two-point estimator that includes the impact of the survey window and scale cuts Then, in Sec. 4, we describe the shear catalog used to test the proposed formalism in this paper and the implementation details of our estimators. The results of the reconstruction from harmonic space and real space with likelihood analysis are shown in Sec. 5. Finally, in Sec. 6, we summarize our results and the future outlook.
@@ 2.1 Cosmic Shear
Cosmic shear is the coherent distortion of images of distant background galaxies by the foreground Large Scale Structure (LSS) in the Universe (see Kilbinger 2015 for a review). Such distortions are induced by all matter along the line of sight, making cosmic shear a powerful tool for probing the matter distribution, providing us with a unique glimpse into the structure and evolution of the Universe on large scales. The first cosmic shear detections were made in the early 2000s (e.g. Kaiser et al., 2000; Bacon et al., 2000; Wittman et al., 2000; Van Waerbeke et al., 2001; Rhodes et al., 2004), and further advancements were made in the subsequent years until now. These improvements included larger volumes of survey data, improved redshift estimation (e.g., Myles et al., 2021; Rau et al., 2023), and refined statistical analysis (e.g., Schrabback et al., 2010). Recent surveys, as well as ongoing ones, have successfully measured cosmic shear and constrained cosmological parameters to high precision. The upcoming imaging surveys like LSST (Ivezić et al., 2019), Euclid (Laureijs et al., 2011), and Nancy Grace Roman Space Telescope (Akeson et al., 2019) will allow us to make percent-level measurements with unprecedented data volume, elevating the importance of the consistency between analyses between the real and harmonic space.
@@ 2.3 Cosmic Shear Analysis
The outcome of the measurements described in Sec. 2.2 is a shear catalog. The shear catalog is then compressed using a summary statistic. Two common summary statistics are to measure the shear-shear power spectrum $C_{\ell}$ in harmonic space and the two-point correlation functions (TPCFs) $\xi_{\pm}(\theta)$ in real space. Higher-order statistics can also be used to extract the non-Gaussian nature of the cosmic density field not encoded by the two-point summary statistics (e.g., Zaldarriaga & Scoccimarro, 2003). Both spaces have their own advantages and disadvantages. It is often easier to estimate the TPCFs (Landy & Szalay, 1993) as our data is in real space, and we have a better understanding of observational systematics in real space. In contrast, Gaussian and non-Gaussian modes and linear and non-linear scales are clearly separated in harmonic space (Hamimeche & Lewis, 2008; Alonso et al., 2019). In harmonic space analysis, individual modes are significantly less correlated compared to real space analysis, making the covariance matrix estimation easier. Our theoretical model of the summary statistics is also computed in harmonic space. For this reason, there has been a large body of literature aimed at developing efficient power spectrum estimators, which have been widely employed in the Cosmic Microwave Background (CMB) two-point measurements (e.g., Wandelt et al., 2001; Brown et al., 2005; Lewis et al., 2001).
The shear field is a spin-2 field, and we can further decompose the lensing signal into the curl-free ($E$ mode) and divergence-free ($B$ mode) components. We use the Limber approximation (Limber, 1953) to compute the model power spectra, which is valid for scales used in this analysis (Dalal et al., 2023). Following this assumption, the cosmic shear power spectra can be related to the matter power spectrum as:
@@
where ${}_{2}d_{\ell,\pm 2}$ is the Wigner-d matrix and we use $H^{\pm}_{\theta,\ell}$ to denote the spherical transform operator The summation is over all $\ell$ and consequently computing $\xi_{\pm}$ to an arbitrary accuracy can be computationally expensive. Here we have added the $B$ mode power spectrum for completeness, and equations (6) and (7) apply to any spin-2 field. Hereafter, we will drop the subscript for tomographic bin pairs and write the multipole modes in the subscript instead.
@@ 3 Method
In this section, we describe the measurement of the cosmic shear power spectrum (Sec. 3.1.2) and two-point correlation functions (TPCFs; Sec. 3.2) from the shear catalog. We discuss the methods to account for survey window and scale cuts, and how to consistently reconstruct the power spectrum from the TPCFs and vice versa in Sec.s 3.3 and 3.4.
@@ 3.1.1 Full Sky Shear Map
where the power spectrum is averaged over the $m$ values for each multipole mode $\ell$. For notational convenience, we assume the two shear fields to be in the same redshift bin and omit explicit expressions for the tomographic redshift bin pairs for the rest of the paper.
@@ 3.1.2 Masked Shear Map
where $\hat{N}_{g}(\bm{n})$ is the observed number of galaxies at position $\bm{n}$. Note that $\bar{N}_{g}$ is a normalization choice we made and as long as normalization choice is treated consistently between data and model (including covariance), it does not have any impact on the analysis (its impact cancels out in the likelihood function). This definition includes clustering and shot noise in the survey window. Note that the denominator in equation (12) is more optimal than dividing by the actual number of galaxies in each pixel as some pixels can have very few galaxies, making the estimation of shear in those pixels noisy. Using a constant in the normalization effectively up-weights the low noise pixels (pixels with more galaxies) and down-weights the high noise pixels (pixels with fewer galaxies). This weighting is included in the model, so the final analysis is unbiased. A more optimal choice of the window includes inverse variance weighting (the FKP weighting), which depends on the signal and noise power spectrum (Feldman et al., 1994; Hamilton, 1997), but for this work, we adopt the weighting scheme of equation (14). In summary, equation (12) is the measured value of shear at pixel $p$, and equation (13) is the theory shear value at pixel $p$ that accounts for the survey window and the pixel window function.
@@
$\hat{D}^{BB}_{\ell}$ $=C^{(\Pi)}_{\ell}\left(M^{+}_{\ell\ell^{\prime}}C^{BB}_{\ell^{\prime}}+M^{-}_{\ell\ell^{\prime}}C^{EE}_{\ell^{\prime}}\right)$ (16)
@@
Additionally, one must account for the noise bias in the estimated power spectrum. The shape noise generates this bias due to the intrinsic ellipticities of galaxies. For a windowed field, the noise is effectively multiplied by the window given by $\sqrt{W^{\gamma}(\bm{n})}$ (Feldman et al., 1994; Li et al., 2019). We analytically estimate the constant shape noise power spectrum as:
@@
where $\sigma_{e}$ is the shape noise and $\bar{n}_{\mathrm{eff}}$ is the effective number of galaxies in inverse steradians. We use the superscript $N$ to denote the coupling matrix for the noise in the shear field as the signal and noise have different windows. We will drop $C^{(\Pi)}_{\ell}$ and $N_{\ell}$ to simplify our notation, though we do correct for the pixel window effect and the additive noise bias on the measurements.

### v2

# Matching Cosmic Shear Analysis in Harmonic and Real Space

@@ Abstract
Recent cosmic shear analyses have exhibited discrepancies of up to $1\sigma$ between the inferred cosmological parameters when analyzing summary statistics in real space versus harmonic space. In this paper, we demonstrate the consistent measurement and analysis of cosmic shear two-point functions in harmonic and real space using the $i$Master algorithm. This algorithm provides a unified prescription to model the survey window effects and scale cuts in both real space (due to observational systematics) and harmonic space (due to model limitations), resulting in a matching estimation of the cosmic shear power spectrum from both harmonic and real space estimators. We show that the $i$Master algorithm gives matching results using measurements from the HSC Y1 mock shape catalogs in both real and harmonic space, resulting in matching inferences of $S_{8}=\sigma_{8}(\Omega_{m}/0.3)^{0.5}$. This method provides an unbiased estimate of the cosmic shear power spectrum, and $S_{8}$ inference that has a correlation coefficient of 0.997 between analyses using measurements in real space and harmonic space when $S_{8}$ is the only free parameter. We observe the mean difference between the two inferred $S_{8}$ values to be 0.0004 across noise-free mock realizations, far below the observed difference of 0.042 for the published HSC Y1 analyses and well below the statistical uncertainties. While the notation employed in this paper is specific to photometric galaxy surveys, the methods are equally applicable and can be extended to spectroscopic galaxy surveys, intensity mapping, and CMB surveys.
@@
pubyear: 2024pagerange: Matching Cosmic Shear Analysis in Harmonic and Real SpaceB
@@ 1 Introduction
Once the data is acquired and processed into shear maps or shear catalogs, the next step is to compute the two-point summary statistics. One question that we need to address at this stage is whether to carry out the cosmic shear analysis in real space or harmonic space; each space has its own advantages and disadvantages, and we introduce the trade-offs in Sec. 2.3. Analysis in real space corresponds to measuring the two-point correlation functions of the galaxy shear estimates as a function of angular separation (e.g., as in Li et al., 2023; Secco et al., 2022; Schneider et al., 2022), and analysis in harmonic space corresponds to measuring the cosmic shear power spectrum as a function of multipole moments (e.g., as in Dalal et al., 2023; Doux et al., 2022; Loureiro et al., 2022). One major challenge is to account for the survey window. The survey window is a function that captures the effects of the survey geometry and other observational effects that affect the selection of galaxies included in our samples. The survey window complicates the power spectrum analysis as the survey mask couples Fourier modes, which would otherwise be independent. This leads to mode coupling in the power spectrum, which is commonly addressed using pseudo-$C_{\ell}$ estimators (e.g., Wandelt et al., 2001; Hikage et al., 2011; Alonso et al., 2019).
@@
Since the coherent distortion of galaxy shapes is caused by a scalar gravitational potential of the matter distribution, we can further separate the lensing signal into curl-free ($E$ mode) and divergence-free ($B$ mode) components. The survey window also introduces leakage of $E$ modes into the $B$ mode component of a given map, and vice versa, an effect that needs to be modeled and corrected using detailed knowledge of the window. The modeling of the survey window in harmonic space was first addressed in Hivon et al. (2002), who introduced the Master algorithm to deconvolve the effects of the survey window to reconstruct the power spectrum of the underlying fields for Cosmological Background Microwave (CMB) analysis (Wandelt et al., 2001; Brown et al., 2005; Lewis et al., 2001). Alonso et al. (2019) introduced the now standard NaMaster software with an implementation of the Master algorithm for spin-0 and spin-2 fields that is also used for galaxy clustering and cosmic shear analyses.
Throughout this paper, we define two estimators as matching if they yield results that agree within a percent when applied to the same dataset, and use the term biased to describe cases where systematic effects cause a significant deviation from the expected result. For example, Hamana et al. (2020) and Doux et al. (2021) show how such a disparity in approach produces an apparent mismatch between the analyses in real space and harmonic space when performing cosmological parameter inference. Crucially, if we truly believe certain scales cannot be well modeled, they should be removed regardless of the space in which the analysis is performed.
The survey window also affects the two-point correlation functions in real space, although this effect is typically ignored since the commonly-used estimator removes most of the window effects (see appendix D of Singh et al. (2020) for demonstration of residual percent-level window effects on correlation functions).
After measuring the summary statistics, we apply scale cuts on the measurements to restrict our analysis to scales where the model is reliable (normally done in harmonic space where the models are defined) and to remove scales contaminated by observational systematics (normally done in real space to account for physical and observational effects tied to angular separation) before performing cosmological parameter inference.
@@
We can typically model a limited range of scales that is smaller than the range of scales probed by the current data. Thus, we need to apply scale cuts carefully to extract as much information as possible while avoiding biases from applying the model outside its range of validity. Failure to account for the scale cut that was implemented in one space when modeling the other space can lead to using different sets of information in real versus harmonic space analysis, even though the underlying field is the same.
@@
There have been many efforts to introduce a statistic that can separate out the $E/B$ modes. Schneider et al. (2002) introduced the band power spectra that used the aperture mass dispersion $\langle M^{2}_{\text{ap}}(\theta)\rangle$ and $\langle M^{2}_{\times}(\theta)\rangle$ (Schneider et al., 1998) measured from a linear combination of the two-point correlation functions weighted by specific window functions to perform the $E/B$ mode decomposition. However, calculating the aperture mass dispersion requires estimation of the shear correlation functions down to zero angular separation, which is not possible in practice. To remedy this, they applied apodization in the window function that truncates the integral over all angular distances and reduces the ringing effect caused by the limited range of the two-point correlation functions.
Schneider et al. (2010) introduced the Complete Orthogonal Sets of E- / B-Mode Integrals (COSEBI) method to measure pure $E$ and $B$-mode cosmic shear power spectra from the two-point correlation functions over a finite interval $\theta_{\text{min}}\leq\theta\leq\theta_{\text{max}}$. The method, similar to Schneider et al. (2002), uses a linear combination of binned measured two-point correlation functions weighted by specific window functions in the angular scale. Becker & Rozo (2016) introduced a similar band-power estimator for cosmic shear power spectra from linear combinations of binned two-point correlation functions. Although the aforementioned methods return unbiased two-point statistics estimators, there were still inconsistencies in analysis between the two spaces.
[…]
@@
Singh (2021) further improved the Master algorithm (to produce $i$Master) and generalized the method to account for the scale cuts used in the analysis, after which the power spectra and two-point correlation function estimates match. In this work, we present the first application of the $i$Master algorithm to survey mock catalogs with complex geometry.
We measure the cosmic shear two-point functions in harmonic and real space in the HSC Y1 mock shape catalogs (Shirasaki et al., 2019), using the $i$Master method to estimate the power spectrum from both harmonic and real space estimators consistently, accounting for both the survey window and the scale cuts used in the analysis in a way that ensures matching results. We quantify the consistency between the two analyses by measuring $\delta S_{8}=S_{8|R\to F}-S_{8|F}$, where $S_{8|R\to F}$ is the inferred $S_{8}$ value going from the real space estimator to the harmonic space estimator, and $S_{8|F}$ using the harmonic space estimator directly.
@@
This paper is organized as follows: In Sec. 2, we provide an overview of cosmic shear analysis. In Sec. 3, we describe how two-point estimators are measured in harmonic and real space and outline steps to make a matching two-point estimator that includes the impact of the survey window and scale cuts. Then, in Sec. 4, we describe the shear catalog used to test the proposed formalism in this paper and the implementation details of our estimators. The results of the reconstruction from harmonic space and real space with likelihood analysis are shown in Sec. 5. Finally, in Sec. 6, we summarize our results and the future outlook.
@@ 2.1 Cosmic Shear
Cosmic shear is the coherent distortion of images of distant background galaxies by the foreground Large Scale Structure (LSS) in the Universe (see Kilbinger 2015 for a review). Such distortions are induced by all matter along the line of sight, making cosmic shear a powerful tool for probing the matter distribution, providing us with a unique glimpse into the structure and evolution of the Universe on large scales. The first cosmic shear detections were made in the early 2000s (e.g. Kaiser et al., 2000; Bacon et al., 2000; Wittman et al., 2000; Van Waerbeke et al., 2001; Rhodes et al., 2004), and further advancements were made in the subsequent years until now. These improvements included larger volumes of survey data, improved redshift estimation (e.g., Myles et al., 2021; Rau et al., 2023), and refined statistical analysis (e.g., Schrabback et al., 2010). Recent surveys, as well as ongoing ones, have successfully measured cosmic shear and constrained cosmological parameters to high precision. The upcoming imaging surveys like LSST (Ivezić et al., 2019), Euclid (Laureijs et al., 2011), and Nancy Grace Roman Space Telescope (Akeson et al., 2019) will allow us to make percent-level measurements with unprecedented data volume. As these datasets become increasingly precise, ensuring matching results between real-space and harmonic-space analyses is essential to prevent systematic biases in cosmological constraints.
@@ 2.3 Cosmic Shear Analysis
The outcome of the measurements described in Sec. 2.2 is a shear catalog. The shear catalog is then compressed using a summary statistic. Two common summary statistics are to measure the shear-shear power spectrum $C_{\ell}$ in harmonic space and the two-point correlation functions (TPCFs) $\xi_{\pm}(\theta)$ in real space. Higher-order statistics can also be used to extract the non-Gaussian nature of the cosmic density field not encoded by the two-point summary statistics (e.g., Zaldarriaga & Scoccimarro, 2003). Both spaces have their own advantages and disadvantages. It is often easier to estimate the TPCFs (Landy & Szalay, 1993) as our data is in real space, and we have a better understanding of observational systematics in real space. In contrast, Gaussian and non-Gaussian modes and linear and non-linear scales are clearly separated in harmonic space (Hamimeche & Lewis, 2008; Alonso et al., 2019). In harmonic space analysis, individual modes are significantly less correlated compared to real space analysis, making the covariance matrix estimation easier. Our theoretical model of the summary statistics is also computed in harmonic space. For this reason, there has been a large body of literature aimed at developing efficient power spectrum estimators, which have been widely employed in the Cosmic Microwave Background (CMB) two-point measurements (e.g., Wandelt et al., 2001; Brown et al., 2005; Lewis et al., 2001). While both approaches should ideally yield matching cosmological constraints, previous studies have found discrepancies between real space and harmonic space analyses when applying different scale cuts or improperly handling the survey function (Hamana et al., 2020; Doux et al., 2021). These mismatches highlight the importance of developing methods that correctly account for scale-dependent effects to avoid introducing systematic biases into cosmological inference.
The shear field is a spin-2 field, and we can further decompose the lensing signal into the curl-free ($E$ mode) and divergence-free ($B$ mode) components. We use the Limber approximation (Limber, 1953) to compute the model power spectra, which is valid for scales used in this analysis (Dalal et al., 2023).
Following this assumption, the cosmic shear power spectra for flat spatial geometry can be related to the matter power spectrum as:
@@
where ${}_{2}d_{\ell,\pm 2}$ is the Wigner-d matrix and we use $H^{\pm}_{\theta,\ell}$ to denote the spherical transform operator. The summation is over all $\ell$ and consequently computing $\xi_{\pm}$ to an arbitrary accuracy can be computationally expensive. Here we have added the $B$ mode power spectrum for completeness, and equations (6) and (7) apply to any spin-2 field. Hereafter, we will drop the subscript for tomographic bin pairs and write the multipole modes in the subscript instead.
@@ 3 Method
In this section, we describe the measurement of the cosmic shear power spectrum (Sec. 3.1.2) and two-point correlation functions (TPCFs; Sec. 3.2) from the shear catalog. We discuss the methods to account for survey window and scale cuts, and how to consistently reconstruct the power spectrum from the TPCFs and vice versa in Sections 3.3 and 3.4. The motivation behind our method is to ensure matching cosmic shear measurements between real-space and harmonic space analyses. Since each space has distinct advantages and challenges, properly accounting for survey masks, pixelization effects, and scale cuts is essential to avoid introducing systematic biases into cosmological inferences.
@@ 3.1.1 Full Sky Shear Map
where the power spectrum is averaged over the $m$ values for each multipole mode $\ell$. For notational convenience, we assume the two shear fields to be in the same redshift bin and omit explicit expressions for the tomographic redshift bin pairs for the rest of the paper. In the idealized case of a full sky, uniform survey, the power spectrum estimators are matching between real and harmonic space because there is no survey mask or scale-dependent selection effects. However, real observations introduce non-trivial window functions and observational systematics that must be accounted for to ensure unbiased results.
@@ 3.1.2 Masked Shear Map
where $\hat{N}_{g}(\bm{n})$ is the observed number of galaxies at position $\bm{n}$. Note that $\bar{N}_{g}$ is a normalization choice we made and as long as normalization choice is treated consistently between data and model (including covariance), it does not have any impact on the analysis (its impact cancels out in the likelihood function). This definition includes clustering and shot noise in the survey window. Using a constant $\bar{N}_{g}$ in the denominator of equation (12) instead of the actual number of galaxies per pixel provides a more stable estimator of the shear. Dividing by the observed number of galaxies in each pixel would increases the noise in low-density pixels, as these estimators are highly sensitive to stochastic fluctuations in regions with fewer galaxies. Instead, the constant $\bar{N}_{g}$ acts as an implicit weighting factor; it reduces the relative contribution of noisy, low-density pixels while maintaining the contribution from high-density, low-noise pixels. Since this weighting is consistently included in the model, the final analysis remains unbiased. A more optimal choice of the window includes inverse variance weighting (the FKP weighting), which depends on the signal and noise power spectrum (Feldman et al., 1994; Hamilton, 1997), but for this work, we adopt the weighting scheme of equation (14). In summary, equation (12) is the measured value of shear at pixel $p$, and equation (13) is the theory shear value at pixel $p$ that accounts for the survey window and the pixel window function.
@@
$\hat{D}^{BB}_{\ell}$ $=C^{(\Pi)}_{\ell}\left(M^{+}_{\ell\ell^{\prime}}C^{BB}_{\ell^{\prime}}+M^{-}_{\ell\ell^{\prime}}C^{EE}_{\ell^{\prime}}\right),$ (16)
@@
Additionally, one must account for the noise bias in the estimated power spectrum. The shape noise generates this bias due to the intrinsic ellipticities of galaxies. For a windowed field, the noise is effectively multiplied by the window given by $\sqrt{W^{\gamma}(\bm{n})}$ (see appendix B of Singh et al. (2020) for derivation of the window and the noise effects in pseudo-$C_{\ell}$ estimators). We analytically estimate the constant shape noise power spectrum as:
@@
where $\sigma_{e}$ is the shape noise, $\bar{n}_{\mathrm{eff}}$ is the effective number of galaxies in inverse steradians, and $M^{N}_{\ell\ell^{\prime}}$ is the coupling matrix from using the window $\sqrt{W^{\gamma}(\bm{n})}$. We use the superscript $N$ to denote the coupling matrix for the noise in the shear field as the signal and noise have different windows. We will drop $C^{(\Pi)}_{\ell}$ and $N_{\ell}$ to simplify our notation, though we do correct for the pixel window effect and the additive noise bias on the measurements.
In summary, this estimator introduces an artificial mixing of $E$- and $B$-modes due to the survey window. If uncorrected, this leads to systematic errors in cosmological parameter estimation.